This problem features the art of finding complex square roots as well as solving a quadratic equation using a resulting formulation based on completing the quadratic square.



(a) Given that u²=-60+32i , express u in the form a+bi where a,b=R



(b) Hence, solve the equation z²-(3_2i)z+5-5i=0

1
Expert's answer
2021-11-07T17:56:21-0500

(a) Let "u=a+bi." Then


"u^2=(a+bi)^2=a^2-b^2+2abi"

Given "u^2=-60+32i." Substitute


"a^2-b^2+2abi=-60+32i"

"\begin{matrix}
 a^2-b^2=-60 \\
 ab=16
\end{matrix}"

"\begin{matrix}
 a^2-\dfrac{256}{a^2}=-60 \\
 b=\dfrac{16}{a}
\end{matrix}"


"a^4+60a^2-256=0"

"D=(60)^2-4(1)(-256)=4624"

"a^2=\dfrac{-60\pm\sqrt{4624}}{2(1)}=-30\pm34"

Since "a\in \R," we take "a^2=-30+34=4"


"\begin{matrix}
 a=-2 \\
 b=-8\end{matrix}"

Or


"\begin{matrix}
 a=2 \\
 b=8\end{matrix}"


"u=-2-8i\ or\ u=2+8i"

b)

"z\u00b2-(3-2i)z+5-5i=0"

"D=(3-2i)^2-4(1)(5-5i)=9-4-12i-20+20i"

"=9-4-12i-20+20i=-15+8i"

"z=\dfrac{-(3-2i)\pm\sqrt{D}}{2(1)}"

Let "u=a+bi." Then


"u^2=(a+bi)^2=a^2-b^2+2abi"

Suppose that "u^2=D=-15+8i." Substitute


"a^2-b^2+2abi=-15+8i"




"\begin{matrix}
 a^2-b^2=-15 \\
 ab=4
\end{matrix}"

"\begin{matrix}
 a^2-\dfrac{16}{a^2}=-15 \\
 b=\dfrac{4}{a}
\end{matrix}"


"a^4+15a^2-16=0"

"D=(15)^2-4(1)(-16)=289"

"a^2=\dfrac{-15\pm\sqrt{289}}{2(1)}=\dfrac{-15\pm17}{2}"

Since "a\in \R," we take "a^2=\dfrac{-15+17}{2}=1"


"\begin{matrix}
 a=-1 \\
 b=-4\end{matrix}"

Or


"\begin{matrix}
 a=1 \\
 b=4\end{matrix}"


"u=-1-4i\ or\ u=1+4i"


"z=\dfrac{-(3-2i)\pm\sqrt{(-1-4i)^2}}{2(1)}"

"=\dfrac{-3+2i\pm(-1-4i)}{2}"

"z_1=\dfrac{-3+2i+1+4i}{2}=-1+3i"

"z_2=\dfrac{-3+2i-1-4i}{2}=-2-i"



"z=\dfrac{-(3-2i)\pm\sqrt{(1+4i)^2}}{2(1)}"

"=\dfrac{-3+2i\pm(1+4i)}{2}"

"z_3=\dfrac{-3+2i-1-4i}{2}=-2-i"

"z_2=\dfrac{-3+2i+1+4i}{2}=-1+3i"

"z=-2-i\ or\ z=-1+3i"